OpenStudy (anonymous):

Find alnumbers for which therational exprssion indefined 3^3-4m/m^2-81

OpenStudy (anonymous):

Assuming I'm reading your expression correct, 3^3 - (4m)/(m^2) - 81... ...the only term you have to worry about is the 4m/(m^2) but note that the m in the numerator cancels out one of the power of m in the denominator so you have 3^3 - 4/m - 81 So the only place the expression is undefined is m = 0. Hope that helps!

jhonyy9 (jhonyy9):

begin with denominator equal zero - so because you know that one fraction than will be undefined sure - so than ther will be (m-9)(m+9)=0 - so m-9=0 so m=9 and m+9=0 so m=-9 - than result that for m=+/- 9 this fraction will be undefined sure

jhonyy9 (jhonyy9):

hope that this will be understandably sure for you good luck bye

OpenStudy (anonymous):

^ plugging 9 and -9 into the expression gives real numbers 3^3 - 4(9)/(9^2) - 81 = 9 - 36/81 - 81 = -652/9 3^3 - 4(-9)/(-9^2) - 81 = 9 + 36/81 - 81 = -644/9 The expression is only undefined when m=0.